# How Temporal Unrolling Supports Neural Physics Simulators

### v1

@@ 5.1 Distribution Shift and Long-term Gradients
These findings are not new [3, 56], but confirm that our setup matches previous work. They already motivate recommendation (IV), i.e. that testing unrolled architectures can be performed efficiently on cheaper low-dimensional problems, such as the KS system.
@@
Next, we investigate the effect of backpropagating gradients in the unrolled chain of network and simulator operations. The non-differentiable NOG setup already addresses the data shift problem, as it exposes the learned dynamics’ attractor at training time. The inference errors are visualized in figure 4. On average, training with NOG over ONE training yields an error reduction of 23%, in line with recommendation II. For large architectures, inference accuracy increases and learned and ground truth systems become more alike. In these cases, their attractors are similar, and unrolling is less crucial to expose the learned attractor. However, NOG training still uses crude gradient approximations, which negatively influences training regardless of how accurate the current model is. We believe this is why the NOG variant falls behind for large sizes. While NOG models remain closer to the target than ONE in all other cases, the results likewise show that the differentiable WIG setup further improves the performance. These networks reliably produce the best inference accuracy. Due to the stochastic nature of the non-linear learning processes, outliers exist, such as the $0.5$m NOG model of the KOLM system. Nonetheless, the WIG models consistently perform best and yield an average improvement of 38% over the ONE baseline. This error average behavior matches best-performing networks’ behavior (C).
Figure 4: Inference accuracy measured in $\mathcal{L}_{2}$ for correction setups on KS, KOLM, and WAKE systems; displayed models were trained with ONE(brown), NOG(blue), WIG(orange); across network architectures (graph networks for KS, conv-nets otherwise
), and network sizes WIG has lowest errors; one 32k ONE model diverged in the KOLM case that NOG and WIG kept stable
Figure 4: Inference accuracy measured in $\mathcal{L}_{2}$ for correction setups on KS, KOLM, and WAKE systems; displayed models were trained with ONE(brown), NOG(blue), WIG(orange); across network architectures (graph networks for KS, conv-nets otherwise
Figure 5 depicts the time until de-correlation between inference runs and the ground truth for the KS system. For all model sizes and both network architectures, WIG achieves the longest inference rollouts until the de-correlation threshold is reached. Similarly to the observations on the $\mathcal{L}_{2}$ metric in Figure 4, NOG is the second best option for small and medium-sized networks when it comes to correlation. This observation again translates between the network architectures.
@@
In addition, we studied the time until divergence for the KS system. This metric sets a high threshold of $\mathcal{L}_{2}>500$, which would correspond to a hundred-fold increase in the state of a simulation. This metric measures the time until the solution blows up, not whether there is a particular similarity to the ground truth. As such, a large number of timesteps can be seen as a metric for the autoregressive stability of the network. Figure 6 shows this evaluation for the KS system. The unrolled setups excel at this metric. Unsurprisingly, mitigating the data shift by unrolling the training trajectory stabilizes the inference runs. Through unrolling, the networks were trained on inference-like states. The most stable models were trained with the WIG setup, whose long-term gradients further discourage unphysical outputs that could lead to instabilities in long inference runs. The large networks’ divergence time comes close to the upper threshold we evaluated, i.e. 1000 steps.
@@
To conclude, our results allow for disentangling the influence of data shift and gradient divergence. As all training modalities, from data sets to random seeds, were kept constant, the only difference between NOG and WIG is the full gradient information provided to the latter. As such, we can deduce from our measurements that reducing the data shift contributes to the aforementioned improvement of 23%, while long-term gradient information yields another improvement of 15% (recommendation II). Figure 7 highlights this by depicting the accuracy of the unrolled setups normalized by the respective ONE setup for different model sizes and physical systems. WIG yields the best $L_{2}$ reductions. NOG in certain cases even performs worse with factors larger than one, due to its mismatch between loss landscape and gradient information, but nonetheless outperforms ONE on average. A crucial takeaway is that unrolling, even without differentiable solvers, provides measurable benefits to network training. As such, interfacing existing numerical codes with machine learning presents an attractive option that does not require the re-implementation of established solvers. Our results provide estimates for the expected gains with each of these options.
Figure 7: $\mathcal{L}_{2}$ error of NOG and WIG setups relative to the error achieved by ONE (i.e. ${\mathcal{L}^{\mathrm{NOG}}_{2}}/{\mathcal{L}^{\mathrm{ONE}}_{2}}$ and ${\mathcal{L}^{\mathrm{WIG}}_{2}}/{\mathcal{L}^{\mathrm{ONE}}_{2}}$); WIG training reliably produces more accurate results than NOG and ONE
@@
The effects of unrolling on data shift and gradient divergence depend on the unrolled length $m$, as summarized in Fig. 1. A long horizon diminishes the data shift, as the learned dynamics can be systematically explored by unrolling. At the same time, gradient inaccuracy impairs the learning signal in the NOG case, while exploding gradients can have a similar effect for long backpropagation chains in the WIG setup. We test this hypothesis by varying the unrolled training horizon. The results are visualized in figure 8. For increasing horizons, the inference performance of NOG variants improves until a turning point is reached. After that, NOG performance deteriorates quickly. In the following regime, the stopped gradients cannot map the information gained by further reducing the data shift to an effective parameter update. The quality of the learning signal deteriorates. This is mitigated by allowing gradient backpropagation through the unrolled chain in the WIG setup. Herein, the inference accuracy benefits from even longer unrollings, and only diverges for substantially larger $m$ when instabilities from recurrent evaluations start to distort the direction of learning updates. These empirical observations confirm the theoretical analysis that went into Figure 1: There exist unrolling horizons for which training performance is improved for both NOG and WIG, while this effective horizon is longer for WIG.
@@
As training with long unrolling poses challenges, we found a curriculum-based approach with an incremental increase of the unrolling length $m$ at training time to be essential [56, 30]. Figure 8 shows how long unrollings in the initial training phases can hurt accuracy. Untrained networks are particularly sensitive since autoregressive applications can quickly lead to divergence of the simulation state or the propagated gradients. At the same time, learning rate scheduling is necessary to stabilize gradients (recommendation IV).
@@
Cutting long chains of gradients was previously proposed as a remedy for training instabilities of unrolling [37, 7, 54]. Discarding gradients for a number of initial steps [7, 47] is indicated by the parameter $w$ describing the number of warm-up steps for which gradients are stopped.
Figure 8: Left: variation of unrolling horizons $m$ trained for the KS system; all models were initialized with the leftmost ONE setup; with NOG training, the error increases and explodes for $m>6$; WIG remains stable for longer horizons, diverges at $m>18$; Middle: different curriculums for the KOLM system; starting the training with large $m$ leads to unfavorable network states; Right: Inference accuracy without learning rate schedule relative to with scheduling, i.e. $\mathcal{L}_{2}^{\mathrm{no-schedule}}/\mathcal{L}_{2}^{\mathrm{schedule}}$; All models benefit from learning rate scheduling, indicated by values are larger than one, and especially for larger models the schedule is crucial
Figure 9: Gradient stopping techniques
As shown in Fig. 9, setting $w$=1 can yield mild improvements over training over the full chain if no curriculum is used. Dividing the backpropagation into subsections [37] likewise does not yield real improvements in our evaluation. We divided the gradient chain into two or three subsections, identified by the parameter $v$ in Fig. 9.
The best performance is obtained with the full WIG setup and curriculum learning, where unrolled models are pre-trained with $m$=1 models. This can be attributed to the more accurate gradients of WIG, as all gradient-stopping variants above inevitably yield a mismatch between loss landscape and learning updates (recommendation II).
@@
Figure 4 shows clear, continuous improvements in accuracy for increasing network sizes. An obvious conclusion is that larger networks achieve better results. However, in the context of scientific computing, arbitrarily increasing the network size is not admissible. As neural approaches compete with established numerical methods, applying pure neural or hybrid architectures always entails accuracy, efficiency, and scaling considerations. The scaling of networks towards large engineering problems on physical systems has been an open question. Figure 10 estimates the convergence rate of the correction networks with respect to the parameter count to be $n^{-{1}/{3}}$. Interestingly, the measured convergence rates are agnostic to the physical system and the studied network architectures. This convergence rate is poor compared to classic numerical solvers, and indicates that neural networks are best applied for their intrinsic benefits. They possess appealing characteristics like data-driven fitting, reduced modeling biases, and flexible applications. In contrast, scaling to larger problems is more efficiently achieved by numerical approaches. In applications, it is thus advisable to combine both methods to render the benefits of both components. In this light, our correction results for small and medium-sized networks have more relevance to potential applications. Therefore, the two main conclusions from this subchapter are as follows: Firstly, the solver part in hybrid architectures scales more efficiently than the neural network part. Secondly, it becomes apparent how network size dominates the accuracy metrics (recommendation III). Thus, parameter counts need to be equal when fairly comparing network architectures. Another observation is the lack of overfitting: despite the largest KS models being heavily overparameterized with five million parameters for a simulation with 48 degrees of freedom accuracy still improves. This finding aligns with recent insights on the behavior of modern large network architectures [5].
@@ 5.2 Varied Learning Tasks
Figure 11: Prediction setups with inference accuracy measured in terms of $\mathcal{L}_{2}$ on KS and WAKE; NOG performs well for small network sizes, while WIG is show advantages for larger ones
@@
The parameter count of models still dominates the accuracy but in contrast to before, the NOG setup performs better than both alternatives for smaller network sizes. This is shown in Fig. 11. The inference errors for the prediction setup are roughly an order of magnitude larger than the respective correction errors. This indicates that pure predictions more quickly diverge from the reference trajectory, especially for small network sizes. The deteriorated performance of WIG can be explained by these larger differences between the inferred state and the references, leading to suboptimal gradient directions. For the setups in Figure 11, NOG training achieves improvements of 31%, and WIG improves on this by a further 3.6%. In general, unrolling maintains strong benefits in prediction setups, but differentiability is less benefitial than in correction setups.
@@
While the learning task is typically mandated by the application, our setup allows us to investigate the effects of unrolling for a smooth transition from prediction to increasingly simple correction tasks, in line with Fig. 2. In the pure prediction case, the physical prior does not model any physics, i.e., is an identity operator. We transition away from pure predictions by providing the network with improving inputs by increasing the time step of the reference solver for correction tasks. On the other end of the spectrum the numerical solver computes the full time step, and the neural network now only has the trivial task to provide an identity function. Since the total error naturally decreases with simpler tasks, figure 13 shows the performance normalized by the task difficulty. The positive effects of NOG and especially WIG training carry over across the full range of tasks. Interestingly, the NOG version performs best for very simple tasks on the right sides of each graph. This is most likely caused by a relatively small mismatch of gradients and energy landscape. In our experiments, the task difficulty was changed by artificially varying the prior’s accuracy. This mimics the effects of basing the correction setups on different numerical schemes. Since unrolling manifests a stable performance improvement across all priors, it promises benefits for many correction setups in other applications. The results above indicate that a 5x accuracy boost can be achieved by integrating low-fidelity physics priors in a NOG setup.
Figure 13: Transition from prediction to correction; a smooth transition in modeling difficulty is achieved by varying the timestep of the numerical solver embedded in the correction step $\gamma=\Delta t_{\mathcal{S}}/\Delta\tilde{t}$, with the solver timestep $\Delta t_{\mathcal{S}}$ and the ground truth timestep $\Delta\tilde{t}$
Figure 14: Left: Prediction of the AERO system, absolute errors of ONE, NOG, and WIG; Right: Relative errors with respect to ONE
@@
The AERO dataset comprises aerofoil flows in the transonic regime and features shocks, fast-changing vorticial structures, and diverse samples around the critical $\text{Ma}=0.8$. As such, we deployed a modern attention-based U-Net [44]. As a complex, large-scale test scenario, our learning setups for the AERO system implement the recommendations we derived from the previous results. Figure 14 shows the inference evaluation of our trained models. Once again, unrolling increases accuracy at inference time. On average, unrolling improves the inference loss by up to 32% in the AERO case. The results reflect our recommendations: Network size has the largest impact, and choosing the right size balances performance and accuracy (III). Additionally, our models were trained with a learning rate scheduled curriculum to achieve the best results (V), details of which are found in Appendix B. Training on the AERO system resembles the behavior of lower-dimensional problems, e.g. the KS system (IV). Similar to our previous prediction tasks, long-term gradients are less crucial, but still deliver the best models (II). However, unrolling itself is essential for stable networks (I).

### v2

@@ 4.1 Distribution Shift and Long-term Gradients
These findings are not new [6, 93], but confirm that our setup matches previous work. They already relate to question (IV), i.e. that testing unrolled architectures can be performed efficiently on cheaper low-dimensional problems, such as the KS system. Figure 7 visualizes exemplary inference rollouts of the KOLM system, where differences in the learning systems emerge.
Figure 5: Exemplary inference trajectories for the KS system on extrapolation data for $\mathcal{X}=70.4$; trajectories of CNNs on the left, errors of CNNs and GCN in the middle, and on the right respectively for all training variants
Figure 6: Exemplary inference trajectories for the WAKE system for the ground truth and WIG training over 50 timesteps
Figure 7: Exemplary inference trajectories over long inference horizons for $Re=1000$, comparing vorticity values for the ground truth simulation, and corrective setups trained with ONE, NOG, and WIG; corrective setups used the 1M network initialized with the same seed
@@
Next, we investigate the effect of backpropagating gradients in the unrolled chain of network and simulator operations. The non-differentiable NOG setup already addresses the data shift problem, as it exposes the learned dynamics’ attractor at training time. The inference errors are visualized in Figure 4. The graph-network setup on the KS is equivalent to the one studied in [10, 20], and the 1M model matches the one used by Brandstetter et al. [10] in terms of size and overall architecture. We present correction results in this section, while previous works have used these graph networks for prediction. The WAKE experiments match the architecture of [57] and span the network size use by Um et al. [93], who also studied correction setups. The closest relative to our KOLM setup can be found in the ”learned correction” by Kochkov et al. [37], where a convolutional ResNet with $\sim 0.2$M parameters was trained in a very similar correction setup.
On average, training with NOG over ONE training yields an error reduction of 33%, which answers question (II). For large architectures, inference accuracy increases and learned and ground truth systems become more alike. We believe that in these cases, their attractors are similar, and unrolling is less crucial to expose the learned attractor. However, NOG training still uses crude gradient approximations, which negatively influences training regardless of how accurate the current model is. We believe this is why the NOG variant falls behind for large sizes. While NOG models remain closer to the target than ONE in all other cases, the results likewise show that the differentiable WIG setup further improves the performance. These networks reliably produce the best inference accuracy. Due to the stochastic nature of the non-linear learning processes, outliers exist, such as the $0.5$m NOG model of the KOLM system. Nonetheless, the WIG models consistently perform best and yield an average improvement of 92% over the ONE baseline. The conclusions from the error average can also be drawn from the behavior of the best-performing networks of each training variation (see D).
@@
Long-term Evaluations
We evaluate a cross-correlation metric to measure the similarity of the data in the inferred trajectories to the ground truth data. We first run an autoregressive inference sequence and calculate the correlation as
$C^{t}=\mathrm{corr}(\mathbf{u}^{t},\tilde{\mathbf{u}}^{t})$ (29)
After calculating the correlation of inference frames to the ground truth for each timestep, we measure the time until the two trajectories are decorrelated. For the trajectories to be decorrelated, we define a threshold of $C<0.8$, and measure the timestep until at least 5 frames in the trajectory exceed this threshold, such that the decorrelation time becomes
$T_{\mathrm{decorr}}=\min_{T}\big{\{}T:\sum_{t=0}^{T}\mathbf{1}(C^{t}<0.8)\geq 5\big{\}}.$ (30)
The statistics of the reached step-counts were gathered over all test cases, as for the $\mathcal{L}_{2}$ errors. Figure 8 depicts the time until de-correlation between inference runs and the ground truth for the KS system. For all model sizes and both network architectures, WIG achieves the longest inference rollouts until the de-correlation threshold is reached. Similarly to the observations on the $\mathcal{L}_{2}$ metric in Figure 4, NOG is the second best option for small and medium-sized networks when it comes to correlation. This observation again translates between the network architectures.
In addition, we studied the time until divergence for the KS system. This metric sets a high $\mathcal{L}_{2}$ threshold, which would correspond to a hundred-fold increase in the state of a simulation, and again measures the timesteps until this value is reached. It can be formalized as
$T_{\mathrm{div}}=\min_{T}\big{\{}T:\mathcal{L}_{2}(\mathbf{u}^{t},\tilde{\mathbf{u}}^{t})\geq 500\big{\}}.$ (31)
Note that this metric does not measure whether there is a particular similarity to the ground truth trajectory. Rather, a large number of timesteps can be seen as a metric for the autoregressive stability of the network. Figure 9 shows this evaluation for the KS system. Here, we also show a correction networks with convolutional networks for the correction of the KS system. These are similar to the ones trained in [55], where convolutional networks of size 0.5k and smaller were used in a correction setting. The unrolled setups excel at this metric. Unsurprisingly, mitigating the data shift by unrolling the training trajectory stabilizes the inference runs. Through unrolling, the networks were trained on inference-like states. This also leads to better alignment of the ground truth and the learned attractor. Figure 9 shows how models trained with ONE are most likely to experience instabilities. The fact that the unrolled setups are more stable indicates that the exploration of the learned attractor at training time improves inference performance. The most stable models were trained with the WIG setup, whose long-term gradients further discourage unphysical outputs that could lead to instabilities in long inference runs. The large networks’ divergence time comes close to the upper threshold we evaluated, i.e. 1000 steps.
@@
To conclude, our results allow for disentangling the influence of data shift and gradient divergence. As all training modalities, from data sets to random seeds, were kept constant, the only difference between NOG and WIG is the full gradient information provided to the latter. As such, we can deduce from our measurements that reducing the data shift contributes to the aforementioned improvement of 33%, while long-term gradient information yields another improvement of 15% (question II). Figure 10 highlights this by depicting the accuracy of the unrolled setups normalized by the respective ONE setup for different model sizes and physical systems. WIG yields the best $L_{2}$ reductions. NOG in certain cases even performs worse with factors larger than one, due to its mismatch between loss landscape and gradient information, but nonetheless outperforms ONE on average. A crucial takeaway is that unrolling, even without differentiable solvers, provides measurable benefits to network training. As such, interfacing existing numerical codes with machine learning presents an attractive option that does not require the re-implementation of established solvers. Our results provide estimates for the expected gains with each of these options.
Figure 10: $\mathcal{L}_{\mathrm{rel}}$ error of NOG and WIG setups relative to the error achieved by ONE (i.e. ${\mathcal{L}^{\mathrm{NOG}}_{\mathrm{rel}}}/{\mathcal{L}^{\mathrm{ONE}}_{\mathrm{rel}}}$ and ${\mathcal{L}^{\mathrm{WIG}}_{\mathrm{rel}}}/{\mathcal{L}^{\mathrm{ONE}}_{\mathrm{rel}}}$); WIG training reliably produces more accurate results than NOG and ONE; shading added for better identification of positive/negative factors
@@
Based on the theoretical analysis in Section 2, the learned simulations should perform best for moderate unrolling horizons and deteriorate when these horizons are further increased. As such, the effects of unrolling on data shift and gradient divergence depend on the unrolled length $m$, as summarized in Figure 1. A long horizon diminishes the data shift, as the learned dynamics can be systematically explored by unrolling. At the same time, gradient inaccuracy impairs the learning signal in the NOG case, while exploding gradients can have a similar effect for long backpropagation chains in the WIG setup. We observe these effects in an experimental setup where the unrolled training horizon is varied for networks where training is continued from an identical starting point obtained via pre-training. The results are visualized in Figure 11. For increasing horizons, the inference performance of NOG variants improves until a turning point is reached. After that, NOG performance deteriorates quickly. In the following regime, the stopped gradients cannot map the information gained by further reducing the data shift to an effective parameter update. The quality of the learning signal deteriorates. As outlined in Section 2, this divergence does not only depend on the underlying timescale, but is also influenced by the number of timesteps. With an increasing number of timesteps, more gradients are cut in the NOG system and the gradients become inaccurate. Nevertheless, it is worth mentioning that the divergence occurs at 6 steps, which is roughly equivalent to $t_{c,\text{KS}}/2=12.4/2$. This inaccuracy effect is mitigated by allowing gradient backpropagation through the unrolled chain in the WIG setup. Herein, the inference accuracy benefits from even longer unrollings and only diverges for substantially larger $m$ when instabilities from recurrent evaluations start to distort the direction of learning updates. As the unrolling horizon $m$ significantly surpasses the timescale $t_{c,\text{KS}}=12.4$, the network optimization produces diminishing returns. These empirical observations confirm the theoretical analysis from Section 2.4: From this theoretical perspective, the chaotic timescale gives an estimate for the unrolling horizon for which gradients explode in the target dynamical system. However, in the early stages of the optimization when training losses are still high, the neural simulator is not a good representation of the target dynamics. In this case, the chaotic timescale of the target is not very meaningful. Thus, the training horizons can not only be determined based on the target dynamics, which is also reflected in our study. Here, we recommend using a curriculum instead as described in Section 2.6.
Figure 11: Left: variation of unrolling horizons $m$ trained for the KS system; all models were pre-trained with the leftmost one-step setup marked as init and then finished training in their specific $m$ parameterizations; with NOG training, the error increases and explodes for $m>6$, which corresponds to $\frac{1}{2}t_{c,\text{KS}}$ characteristic timescales in our discretization; WIG remains stable for longer horizons, diverges at $m>18$, i.e. for $m>t_{c,\text{KS}}$; Middle: different curriculums for the KOLM system; starting the training with large $m$ leads to unfavorable network states; Right: Inference accuracy without learning rate schedule relative to with scheduling, i.e. $\mathcal{L}_{2}^{\mathrm{no-schedule}}/\mathcal{L}_{2}^{\mathrm{schedule}}$; All models benefit from learning rate scheduling, indicated by values are larger than one, and especially for larger models the schedule is crucial
@@
The curriculum-based approach features an incremental increase of the unrolling length $m$ at training time [93, 39]. Figure 11 shows how long unrollings in the initial training phases can hurt accuracy. Untrained networks are particularly sensitive since autoregressive applications can quickly lead to divergence of the simulation state or the propagated gradients. At the same time, learning rate scheduling is necessary to stabilize gradients (question IV).
While designing unrolling horizons is not trivial, the fundamental takeaway is that there exist such horizons for which training performance is improved for both NOG and WIG, while this effective horizon is longer for WIG, as illustrated in Figure 1.
@@
Cutting long chains of gradients was previously proposed as a remedy for training instabilities of unrolling [50, 10, 86]. A popular option is to discard gradients for a number of initial steps. This technique is referred to as ”pushforward trick” in [10] and as ”warm-up” steps in [70]. This procedure is similar to the unrolling variants we study. Like unrolling, it time-advances the system, thus drawing data from a more inference-like distribution. However, there are significant differences in the treatment of the losses on intermediate steps and their backpropagation gradients. In the warm-up approach, a number $w$ of initial steps are discarded for loss calculation, and their gradients are stopped. Only steps after the warm-up horizon actually contribute to the loss and are differentiated. The loss formulation is given by
$\mathcal{L}_{w}=\sum_{s=w+1}^{m}\mathcal{L}_{2}(\tilde{u}^{t+s},g_{\theta}^{s}(\tilde{u}^{t}))$ (32)
Another strategy to stabilize unrolling consists of truncating the backpropagation gradients [89]. It was reported to be beneficial in long unrolled physical systems [50]. In this approach, the forward process of the unrolling is not touched at all. However, the backpropagation paths are only evaluated in subsections of the full unrolled trajectory, effectively limiting the maximum length of a backpropagation path. The gradient propagation for unrolling horizons $m$ split into $v$ subsections reads as
$\frac{\partial\mathcal{L}_{\mathrm{v}}^{s}}{\partial\theta}=\sum_{B=B_{v}}^{s}\bigg{[}\frac{\partial\mathcal{L}^{s}}{\partial g_{\theta}^{s}}\bigg{(}\prod_{b=1}^{s-B}\frac{\partial g_{\theta}^{s-b}}{\partial g_{\theta}^{s-b-1}}\bigg{)}\frac{\partial g_{\theta}^{B}}{\partial\theta}\bigg{]},\quad\mathrm{with}\quad B_{v}=\bigg{(}\left\lfloor\frac{s-1}{m/v}\right\rfloor\frac{m}{v}\bigg{)}+1.$ (33)
We evaluate these methods in comparison to unrolled models trained with and without unrolling.
Figure 12: Evaluation of gradient stopping techniques; $m=6$ shows performance of networks trained without curriculum with 6 unrolling steps activated from the start; warm-up steps with $w=1$ discarded steps at the beginning yields minor improvements; backpropagation splitting shows no improvements over the differentiable WIG; best performance is reached with unrolling and curriculum, starting with the init network before unrolling 6 steps
As shown in Figure 12, setting $w$=1 can yield mild improvements over training with the full chain if no curriculum is used, which is in line with the results in [10, 70]. Dividing the backpropagation into subsections [50] likewise does not yield real improvements in our evaluation. We divided the gradient chain into two or three subsections, identified by the parameter $v$ in Figure 12.
The best performance is obtained with the full WIG setup and curriculum learning, where unrolled models are pre-trained with $m$=1 models. This can be attributed to the more accurate gradients of WIG, as all gradient-stopping variants above inevitably yield a mismatch between loss landscape and learning updates (question II).
@@
Figure 4 shows clear, continuous improvements in accuracy for increasing network sizes. An obvious conclusion is that larger networks achieve better results. However, in the context of scientific computing, arbitrarily increasing the network size is not admissible. As neural approaches compete with established numerical methods, applying pure neural or correction approaches always entails accuracy, efficiency, and scaling considerations. The scaling of networks towards large engineering problems on physical systems has been an open question. We compute a convergence rate between the test loss and the number of network parameters. To this end, we use an average test loss for each individual combination of network size and training setup. Here we use the $\mathcal{L}_{2}$ metric following the standard definition of convergence in numerical analysis [34]. This metric provides consistency across timesteps. Relative errors result in unequally weighted timeframes when ground truth solution values vary across time. This is to be avoided for the medium-range test horizons investigated here (40 steps in KS, 250 steps KOLM). The average is computed over the full set of random seeds used in our study, i.e., 8 to 20 individual training runs per variant and size depending on the physical system. The convergence rate estimates are based on the same test trajectories as previous evaluations like Figure 4. For the correction setups, we estimate the convergence rate of the correction networks with respect to the parameter count $n$ to be $n^{-{1}/{3}}$, as shown in Figure 13.
Interestingly, the measured convergence rates persist across the tested physical system and the studied network architectures. This convergence rate is poor compared to classic numerical solvers. Consequently, neural networks are best applied for their intrinsic benefits which derive from the universal approximation theorem [15]. They possess appealing characteristics like data-driven fitting, reduced modeling biases, and flexible applications. In contrast, scaling to larger problems is more efficiently achieved by numerical approaches. In applications, it is thus advisable to combine both methods to render the benefits of both components. In this light, our correction results for small and medium-sized networks have more relevance to potential applications. Therefore, the two main conclusions from this subchapter are as follows: Firstly, the solver part in correction approaches scales more efficiently than the neural network part. Secondly, it becomes apparent how network size dominates the accuracy metrics (question III). Thus, parameter counts need to be equal when fairly comparing network architectures.
@@
Another observation is the lack of overfitting: despite the largest KS models being heavily overparameterized with five million parameters for a simulation with only 48 computational cells accuracy still improves. This finding aligns with recent insights on the behavior of modern large network architectures [8].
@@ 4.2 Varied Learning Tasks
Figure 14: Prediction setups with inference accuracy measured in terms of $\mathcal{L}_{\mathrm{rel}}$ on KS and WAKE; NOG performs well for small network sizes, while WIG is show advantages for larger ones
@@
The parameter count of models still dominates the accuracy but in contrast to before, the NOG setup performs better than both alternatives for smaller network sizes. This is shown in Figure 14. The KS prediction setups are comparable to [84], where convolutional ResNets of size 0.2M were trained on prediction tasks, albeit their networks feature an additional ”dilation” ratio, specifying a sparsely populated filter. The WAKE predictions are similar to [58]. The inference errors for the prediction setup are roughly an order of magnitude larger than the respective correction errors. This indicates that pure predictions more quickly diverge from the reference trajectory, especially for small network sizes. For the setups in Figure 14, NOG training achieves improvements of 41%, and WIG improves on this by a further 30%. In general, unrolling maintains strong benefits in prediction setups, but differentiability is less beneficial than in correction setups. Similarly to the correction task before, we use these test trajectories to estimate a convergence rate with respect to increasing parameter counts. Results are visualized in Figure 15. The prediction convergence rate is slightly worse but comparable to the correction setups.
@@
While the learning task is typically mandated by the application, our setup allows us to investigate the effects of unrolling for a smooth transition from prediction to increasingly simple correction tasks, in line with Figure 2. This is achieved by introducing a parameterized timestep $\Delta t=\gamma\Delta\bar{t}$ to the correction setup, where $\Delta\bar{t}$ represents the ground truth timestep. In contrast to the previous correction setup for the KS equation, we use the ground truth solver $\bar{\mathcal{S}}$ for this setup. Furthermore, as in the regular KS setup, the physical state does not experience any reduction, such that $R(\bar{\mathbf{u}})=\bar{\mathbf{u}}$. The only source of error between the ground truth and the correction setup stems from the different time step sizes and thus depends on the value of $\gamma$. We formalize the setup as
$\mathbf{u}^{t+1}=f_{\theta}(\bar{\mathcal{S}}(\mathbf{u}^{t};\gamma\Delta\bar{t})).$ (34)
With this setup a prediction task without any numerical solver is constructed by setting $\gamma=0$, while $\gamma=1$ recovers the ground truth system where no correction is necessary. It is evident that the latter does not constitute a meaningful learning task, but for all values in between, the difficulty of correction tasks increases with decreasing $\gamma$. In other words, we transition away from pure predictions by providing the network with improving inputs by increasing the time step of the reference solver for correction tasks. Since the total error naturally decreases with simpler tasks, Figure 16 shows the performance normalized by the task difficulty $1-\gamma$. The positive effects of NOG and especially WIG training carry over across the full range of tasks. Interestingly, the NOG version performs best for very simple tasks on the right sides of each graph. This is most likely caused by a relatively small mismatch of gradients and energy landscape. As the task difficulty was changed by artificially varying the numerical solver’s accuracy, this mimics the effects of basing the correction setups on different numerical schemes. Unrolling manifests a stable performance improvement across all numerical accuracies, from very crude approximations (10%) to highly accurate ones that almost fit the ground truth on their own (90%). We also observe these effects of unrolling for different architectures in convolutional and graph-based networks. Thus, unrolling promises benefits for many other correction setups utilizing numerical solvers of with different accuracy and various neural architectures.
Figure 16: Transition from prediction to correction; a smooth transition in modeling difficulty is achieved by varying the timestep of the numerical solver embedded in the correction step $\gamma=\Delta t/\Delta\bar{t}$, with the solver timestep $\Delta t$ and the ground truth timestep $\Delta\bar{t}$; results are normalized by the correction difficulty $1-\gamma$
Figure 17: Left: Prediction of the AERO system, $\mathcal{L}_{\mathrm{rel}}$ errors of ONE, NOG, and WIG; Right: Relative errors with respect to ONE
@@
As a complex, large-scale test scenario, our learning setups for the AERO system implement the recommendations we derived from the previous results. This dataset comprises aerofoil flows in the transonic regime and features shocks, fast-changing vorticial structures, and diverse samples around the critical $\text{Ma}=0.8$. As such, we deploy a modern attention-based U-Net [62]. Figure 17 shows the inference evaluation of our trained models. Once again, unrolling increases accuracy at inference time. On average, unrolling improves the inference loss by up to 41% in the AERO case. The results reflect our research targets: Network size has the largest impact, and choosing the right size balances performance and accuracy (III). Additionally, our models were trained with a learning rate scheduled curriculum to achieve the best results, details of which are found in C. Training on the AERO system resembles the behavior of lower-dimensional problems, e.g. the KS system (IV). Similar to our previous prediction tasks, long-term gradients are less crucial but still deliver the best models (II). However, unrolling itself is essential for stable networks (I).
Extension to three dimensions
To demonstrate that unrolling is effective regardless of the system’s dimensionality, we conduct experiments on the three-dimensional KS equation. We trained a convolutional ResNet to correct a lower-order solver, details as previously described in Section 3.4. The results are visualized in Figure 18. A familiar pattern emerges as unrolled variants outperform the ONE baseline. Once again, fully differentiable training with WIG performs best, while non-differentiable NOG unrolling comes in second. While this test series focuses on a single, fixed network size, we can still observe that unrolling behaves similarly to the previous cases in higher dimensions.
Figure 18: Comparison of the training variants on the three-dimensional KS system, set up as a correction tasks with a convolutional ResNet; $\mathcal{L}_{\mathrm{rel}}$ evaluation on the left; visualization of the system state after 20 steps of simulation on the right